Operations Report – February 14th

Crew IKEA Operations Report 14 Feb 2018

SOL: 4

Name of person filing report: Shannon Rupert

Non-nominal systems: Power System Control

Notes on non-nominal systems: Nothing new to report

Generator (hours run): Due to clouds and contractors, the generator was run all day. The guys needed power for welding and sanding, so even when the sun did come out, we decided to just keep it on.

Solar— 99 % SOC

Diesel – 72 %

Propane – 48 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 0 gallons.

Water (trailer) – 0 gallons.

Water (static) – 450 gallons (added 80 gallons)

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: not recorded

Toilet tank emptied: no

ATV’s Used: yes

Oil Added? n/a

ATV Fuel Used: n/a Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: I started and warmed them up today.

Deimos rover used: Yes

Hours: moved to another parking space

Beginning charge:

Ending charge:

Currently charging: No

Sojourner rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: Yes

Hours: Moved to another parking space

Beginning charge:

Ending charge:

Currently charging: No

Opportunity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

Curiosity rover used: Yes

Hours: Moved to another parking space

Beginning charge:

Ending charge:

Currently charging: No

HabCar used and why, where? Yes, moved to Hab to pump the remaining water into the static tank. My truck is dead (apparently the battery has gone bad as there is battery acid on top) and so tomorrow we will use the HabCar to go to town and get a new battery for the truck and also pick up the packages that are in town.

General notes and comments: Still muddy most of the day but a strong wind later in the afternoon has dried things out quickly.

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: The new stairs and ladders arrived! The contractors began installation and one of the two ladders has been put into place. They are currently bare metal only. We will need to decide which looks best, polished steel or paint. It looks like we will not have time to finish them before the crew arrives on Saturday, so we won’t have to decide right away. Finishing them will probably happen in early May.

Summary of GreenHab operations: Atila continued working on the aquaponics set up and the reorganization of plants. We harvested some greens for salads tonight, and eyed the cucumbers. They may not last until the crew arrives! In addition, we doubled up the pots for the plants in an effort to maintain moisture in the soil longer. In the future, potting soil should be mixed with perlite to ameliorate this problem.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing at this time

Sol Summary – February 14th

Crew IKEA Sol Summary 14 Feb 2018

SOL 4:

Summary Title: Here comes the sun
Author’s name: Atila Meszaros

Mission Status: All nominal

Sol Activity Summary:

Busy day for everyone. The contractor came and installed the new ladders for the upper deck and just before the end of the week the new lower deck ladder is going to be installed. They look fantastic! I’ll love to be here whit the Ikea project will be complete.

Sadly, the mud persisted due to the lack of wind and warmer weather, so we couldn’t get out and do some field work, tomorrow I bet will be luckier. Instead of going out, I worked all day on the Green Hab on the aquaponics, doing some harvest and organizing everything.

Everything is looking good on Mars, but the wind is sometimes scary, inside the Green Hab, you can almost feel the wind screaming. It’s pretty amazing, though.

Thank you for being with us tonight, Sylvain!

Atila

Look Ahead Plan:

Getting a new battery for Shannon’s car, taking Rovers out and I’m rooting for finally some field work and exploration!

Anomalies in work:

Installation of brand new ladders for the upper deck
Shannon’s car battery died

Weather: Temperature: 6°C

Crew Physical Status: Healthy

Reports to be filed:

Sol Summary
Operations Report

Support Requested: None